OpenStudy (anonymous):

@callisto, help here please:) Help here please: Find the exact value of the following, show all work: a. sin^2(75°) / sin 45° cos^2 30° (Hint: sin^2 θ = (sin θ)^2):

OpenStudy (callisto):

Not an easy one ... cos2θ = 2cos^θ-1 Now put θ=15° cos 2(15°) = 2cos^2 15° -1 cos 30 = 2cos^ 15° -1 sqrt(3) /2 = 2cos^2 15° -1 (sqrt3 + 2) /2 = 2cos^215° cos15° = sqrt{(sqrt3 + 2)} /2 ------------------------------------------------------------------ \[sin^2(75°) / ( sin 45° cos^2 30° )\]\[= cos ^2(90°-75°)/ ( sin 45° cos^2 30° )\]\[= cos ^2(15°)/ ( sin 45° cos^2 30° )\]\[= [\sqrt{(\sqrt3 + 2)} /2]^2/ [(\sqrt2/2) (\sqrt3 /2 )^2]\]\[= [(\sqrt3 + 2 /4)/ [(\sqrt2/2) (\sqrt3 /2 )^2]\]\[= [(\sqrt3 + 2 /4)/ [(\sqrt2/2) (3 /4 )]\]\[= [(\sqrt3 + 2 /4)/ [(3\sqrt2/8)]\]\[= 2(\sqrt3 + 2 )/ (3\sqrt2)\]\[= 2\sqrt2(\sqrt3 + 2 )/ (3\times2)\]\[= \sqrt2(\sqrt3 + 2 )/ (3)\]\[= (\sqrt6 + 2\sqrt2 )/ (3)\] Sorry , just woke up :(

OpenStudy (anonymous):

Omg, Thank you so much, that must have been miserable to type. Can you give me a couple of minutes to analyse this? Can I ask you questions after a couple of mins?

OpenStudy (callisto):

Welcome...my method is a little....stupid...perhaps..... SORRY!!!!!

Directrix (directrix):

I went this route but stopped before I did the denominator part. I was trying to avoid that half-angle formula. sin(A+B)=sin A cos B + cos A sin B sin(75) = sin(30 + 45) = sin 30 *cos 45 + cos 30* sin 45 sin(75) = (1/2) * sqrt(2)/2 + sqrt(3)/2*sqrt(2)/2 sin (75) = [ sqrt(2) + sqrt(6) /4 ] [sin (75) ] ^2 = [ sqrt(2) + sqrt(6) /4 ]*[ sqrt(2) + sqrt(6) /4 ]

OpenStudy (anonymous):

Yeah, I got that numerator already:)

OpenStudy (anonymous):

weird part is solving the denominator and then simplifying

Directrix (directrix):

@Open2study -> the denominator is 3 sqrt(2)/8, isn't it?

OpenStudy (anonymous):

Yes it is

Directrix (directrix):

I can't read what Callisto got because the translator or whatever is not working.

OpenStudy (callisto):

Sorry, that's why I said my method was... stupid :(

OpenStudy (anonymous):

@callisto, Dont say that! your methods always help me get great answers:)

OpenStudy (callisto):

Not this time though

Directrix (directrix):

@Callisto --> It's not stupid. It's this program that converts MathLab to text or AjaxMath or something.

OpenStudy (anonymous):

when answering this type of questions, do I continue untill i get simplest form?

Directrix (directrix):

Numerator: [ sqrt(2) + sqrt(6) /4 ]*[ sqrt(2) + sqrt(6) /4 ] Denominator: 3 sqrt(2)/8 Trig is always tedious like this.

Directrix (directrix):

So what can we simplify?

OpenStudy (anonymous):

the denominators, 4, 4, and 8

OpenStudy (anonymous):

right?

Directrix (directrix):

The numerator when squared out is: 19/8 + sqrt(3) or [19 + 8 sqrt(3)] / 8

Directrix (directrix):